1. I invested $23,000 exactly 7 years ago and the balance is now $86,000. What effective annualised interest rate did I earn? (as a percentage to the closest two decimal places; do not show the % sign. E.g 2.876% is 2.88).

2 .Jack has accumulated $99,000 from an initial investment of just $17,000. If the investment earned interest at a rate of 12.9% p.a compounded annually how long ago did Jack make the initial investment? Answer in years to two decimal places.
Answer #1

1.
=(86000/23000)^(1/7)-1=20.7325500313327%

2.
=LOG(99000/17000)/LOG(1+12.9%)=14.5213329135414